# Precompactness in matrix weighted Bourgain-Morrey spaces

### v1

@@ 4 Dyadic average operator conditions for precompact sets
$\lim_{a\to-\infty,a\in\mathbb{Z}}\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}=0,$ (4.1)
@@
$E_{d,a}\vec{f}(x):=\sum_{Q\in{\mathcal{D}}_{-a}}\frac{\chi_{Q}(x)}{|Q|}\int_{Q}\vec{f}(y){\mathrm{d}}y.$ (4.2)
Recall that $x_{Q}$ is the left corner of $Q\in{\mathcal{D}}$. Then for $x\in Q\in\mathcal{D}_{-a}$
@@
Then we will prove that (3.1) is stronger than (4.1). Indeed, suppose that condition (3.1) holds. For any $\varepsilon>0$, there exists a cube $R$ with center 00 and side length $02^{a+1}>$ 0 such that
$\sup_{\vec{f}\in\mathcal{F}}\sup_{y\in R}\|\vec{f}-\tau_{y}\vec{f}\|_{M_{p}^{t,r}(W)}<\varepsilon.$ (4.3)
If $x,y\in Q\in{\mathcal{D}}_{-a}$, then $x-y\in R.$ By Jensen’s inequality ($p\geq 1$), Hölder’s inequality ($r/p\geq 1$), (4.3), we obtain
$\|\vec{f}-E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}=\left(\sum_{Q\in{\mathcal{D}}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|\frac{1}{|Q_{-a,0}|}\right.\right.\right.$
×$\quad\times\left.\left.\left.\int_{Q_{-a,0}}W^{1/p}(x)\left(\vec{f}(x)-\chi_{Q_{-a,0}}(x-x_{Q})\vec{f}(y-x_{Q})\right){\mathrm{d}}y\right|^{p}{\mathrm{d}}x\right)^{r/p}\right)^{1/r}$
$\leq\left(\sum_{Q\in{\mathcal{D}}}W(Q)^{r/t-r/p}\left(\frac{1}{|Q_{-a,0}|}\int_{Q_{-a,0}}\right.\right.$
×$\quad\times\left.\left.\int_{Q}\left|W^{1/p}(x)\left(\vec{f}(x)-\chi_{Q_{-a,0}}(x-x_{Q})\vec{f}(y-x_{Q})\right)\right|^{p}{\mathrm{d}}x{\mathrm{d}}y\right)^{r/p}\right)^{1/r}$
$\leq\left(\sum_{Q\in{\mathcal{D}}}W(Q)^{r/t-r/p}\frac{1}{|Q_{-a,0}|}\int_{Q_{-a,0}}\right.$
×$\quad\times\left.\left(\int_{Q}\left|W^{1/p}(x)\left(\vec{f}(x)-\chi_{Q_{-a,0}}(x-x_{Q})\vec{f}(y-x_{Q})\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}{\mathrm{d}}y\right)^{1/r}$
$\leq\frac{1}{|Q_{-a,0}|^{1/r}}|Q_{-a,0}|^{1/r}\sup_{y\in R}\left(\sum_{Q\in{\mathcal{D}}}W(Q)^{r/t-r/p}\right.$
×$\quad\times\left.\left(\int_{Q}\left|W^{1/p}(x)(\vec{f}(x)-\vec{f}(x-y))\right|^{p}{\mathrm{d}}x\right)^{r/p}\right)^{1/r}$
$\leq\varepsilon,$
modified when $r=\infty$. Hence we prove that (3.1) is stronger than (4.1).
@@
where $E_{d,a}$ is same as (4.2).
  • Proof
    Assume that $\mathcal{F}\subset M_{p}^{t,r}(W)$ satisfies (i)-(iii). Given $\varepsilon>0$ small enough, to prove the total boundedness of $\mathcal{F}$, it suffices to find a finite $\varepsilon$-net of $\mathcal{F}$. Denote by $R_{i}:=[-2^{i},2^{i})^{n}$ for $i\in\mathbb{Z}$. Then from condition (ii), there exists a positive integer $m$ large enough such that
    $\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-\vec{f}\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}<\varepsilon.$ (4.4)
@@
    $\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}<\varepsilon.$ (4.5)
    There exists a sequence $\{Q_{j}\}_{j=1}^{N}$ of disjoint cubes in ${\mathcal{D}}_{-a}$ such that $R_{m}=\cup_{i=1}^{N}Q_{j}$, where $N=2^{(m+1-a)n}$.
    For any $\vec{f}\in\mathcal{F}$ and $x\in{\mathbb{R}^{n}}$, let
    $\varPhi(\vec{f})(x):=\begin{cases}\vec{f}_{Q_{j}}:=\frac{1}{|Q_{j}|}\int_{Q_{j}}\vec{f}(y){\mathrm{d}}y,&x\in Q_{j},j=1,2,\ldots,N,\\ \vec{0},&\operatorname{otherwise}.\end{cases}$
@@
    $\varPhi(\vec{f})(x)=E_{d,a}\vec{f}(x).$
    Hence via (4.4) and (4.5), we have
    $\|\vec{f}-\varPhi(\vec{f})\|_{M_{p}^{t,r}(W)}$ $\leq\|(\vec{f}-\varPhi(\vec{f}))\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}+\|(\vec{f}-\varPhi(\vec{f}))\chi_{R_{m}^{c}}\|_{M_{p}^{t,r}(W)}$ (4.6)
    $=\|(\vec{f}-E_{d,a}\vec{f})\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}+\|\vec{f}\chi_{R_{m}^{c}}\|_{M_{p}^{t,r}(W)}$ (4.7)
    $\leq\|(\vec{f}-E_{d,a}\vec{f})\|_{M_{p}^{t,r}(W)}+\varepsilon$ (4.8)
    $\leq 2\varepsilon.$ (4.9)
    From (4.9), it suffices to show that $\varPhi(\mathcal{F})$ is totally bounded in $M_{p}^{t,r}(W)$. And we have proved that $\varPhi(\mathcal{F})$ is totally bounded in $M_{p}^{t,r}(W)$ in Theorem 3.2. Thus the proof of Theorem 4.1 is complete. $\Box$
Lemma 4.2 (Lemma 4.5, [8])
@@
  • Proof
    First, for any $\vec{f}=(f_{1},f_{2},\ldots,f_{d})^{T}\in M_{p}^{t,r}(W)$, it suffices to show that $f_{i}\in L^{1}_{\operatorname{loc}}({\mathbb{R}^{n}})$ for each $i=1,2,\ldots,d.$ Since $W\in\mathcal{A}_{p}$, then by (2.3)
@@
    $\int_{Q}|\vec{f}|{\mathrm{d}}x$ $=\int_{Q}|\vec{f}|\|W^{-1}\|^{-1/p}\|W^{-1}\|^{1/p}{\mathrm{d}}x$
    $\leq\left(\int_{Q}|\vec{f}|^{p}\|W^{-1}\|^{-1}{\mathrm{d}}x\right)^{1/p}\left(\int_{Q}\|W^{-1}\|^{p^{\prime}/p}{\mathrm{d}}x\right)^{1/p^{\prime}}$
    $\leq C_{Q,W,p}\left(\int_{Q}|W^{1/p}\vec{f}|^{p}{\mathrm{d}}x\right)^{1/p}$
    $\lesssim\|\vec{f}\|_{M_{p}^{t,r}(W)}<\infty.$
@@
Theorem 4.5
Let $W\in\mathcal{A}_{p}$ has the $\mathcal{A}_{p}$-dimension $\tilde{d}\in[0,n)$ with the doubling exponent $\beta$.
(i) Let $1\leq p<t<r<\infty$ If $(\tilde{d}-n+\beta)r/p+n-\beta r/t<0$, then for each $a\in\mathbb{Z}$, $E_{d,a}$ is a bounded operator on matrix weighted Bourgain-Morrey spaces $M_{p}^{t,r}(W)$.
(ii) Let $1\leq p\leq t<r=\infty$ If $\tilde{d}/p-n/p-\beta(1/t-1/p)\leq 0$, then for each $a\in\mathbb{Z}$, $E_{d,a}$ is a bounded operator on matrix weighted Bourgain-Morrey spaces $M_{p}^{t,r}(W)$.
  • Proof
@@
    $+\left(\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}\left(\int_{Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}\right)^{1/r}$
    $=:S_{1}+S_{2}.$
@@
    $S_{1}$ $\leq C_{n,d,p,W}\left(\sum_{j\leq-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}\left(\int_{Q}|W^{1/p}(x)\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}\right)^{1/r}$
    $\lesssim\|\vec{f}\|_{M_{p}^{t,r}(W)}.$
    When $j>-a$, we denote by $(j+a)_{\operatorname{pa}}Q\in{\mathcal{D}}_{-a}$ the $(j+a)$-th dyadic parent of $Q\in{\mathcal{D}}_{j}$. Let $\{A_{Q}\}_{Q\in{\mathcal{D}}}$ be the reducing operator of order $p$ for $W$. By Lemmas 2.8, we have
    $\|A_{Q}A_{(j+a)_{\operatorname{pa}}Q}^{-1}\|\leq c2^{(j+a)\tilde{d}/p}.$ (4.10)
    In $S_{2}$, using (2.5) and (4.10) , we have
@@
    $=\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}$
    ×$\quad\times\left(\frac{|Q|}{|Q|}\int_{Q}\left|W^{1/p}(x)\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}\vec{f}(y){\mathrm{d}}y\right|^{p}{\mathrm{d}}x\right)^{r/p}$
@@
    $\lesssim\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}|Q|^{r/p}$
    ×$\quad\times\left(2^{(j+a)\tilde{d}/p}\left|A_{(j+a)_{\operatorname{pa}}Q}\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}\vec{f}(y){\mathrm{d}}y\right|\right)^{r}$
    $\approx\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}|Q|^{r/p}2^{(j+a)\tilde{d}r/p}$
    ×$\quad\times\left(\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}\vec{f}(y){\mathrm{d}}y|^{p}{\mathrm{d}}x\right)^{r/p}$
    $=\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}$
    ×$\quad\times\left(\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}$
    $=\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}2^{-anr/p}$
    ×$\quad\times\left(\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}.$
    Remark that given $S\in{\mathcal{D}}_{-a}$, there are $2^{(j+a)n}$ cubes $R$ such that $(j+a)R=S$. Hence, by Lemma 4.4 and $(\tilde{d}-n+\beta)r/p+n-\beta r/t<0$, we have
    $\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}2^{-anr/p}$
    ×$\times\left(\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}$
    $=\sum_{j>-a}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}2^{-anr/p}$
    ×$\quad\times\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}\left(\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}$
    $\leq\sum_{j>-a}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}2^{-anr/p}2^{(j+a)n}2^{-(j+a)\beta(r/t-r/p)}$
    ×$\quad\times\sum_{S\in\mathcal{D}_{-a}}W(S)^{r/t-r/p}\left(\int_{S}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}$
    $\lesssim\|\vec{f}\|_{M_{p}^{t,r}(W)}^{r},$
    where in the second step we use the following fact. Since $W((j+a)_{\operatorname{pa}}Q)\leq 2^{(j+a)\beta}W(Q)$ and $01/t-1/p<$ 0, we have
    $W(Q)^{1/t-1/p}\leq 2^{-(j+a)\beta(1/t-1/p)}W((j+a)_{\operatorname{pa}}Q)^{1/t-1/p}.$
@@
$\lesssim W(Q_{j,k})^{1/t-1/p}\left(\int_{Q_{j,k}}|W^{1/p}(x)\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{1/p}.$
If $j>-a$, we denote by $(j+a)_{\operatorname{pa}}Q_{j,k}\in{\mathcal{D}}_{-a}$ the $(j+a)$-th dyadic parent of $Q_{j,k}\in{\mathcal{D}}_{j}$.

### v2

@@ 4 Dyadic average operator conditions for precompact sets
$\lim_{a\to-\infty,a\in\mathbb{Z}}\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}=0,$ (18)
@@
$E_{d,a}\vec{f}(x):=\sum_{Q\in{\mathcal{D}}_{-a}}\frac{\chi_{Q}(x)}{|Q|}\int_{Q}\vec{f}(y){\mathrm{d}}y.$ (19)
Recall that $x_{Q}$ is the lower left corner of $Q\in{\mathcal{D}}$. Then for $x\in Q\in\mathcal{D}_{-a}$
@@
Then we will prove that (8) is stronger than (18). Indeed, suppose that condition (8) holds. For any $\epsilon>0$, there exists a cube $R$ with center 0 and side length $2^{a+1}>0$ such that
$\sup_{\vec{f}\in\mathcal{F}}\sup_{y\in R}\|\vec{f}-\tau_{y}\vec{f}\|_{M_{p}^{t,r}(W)}<\epsilon.$ (20)
If $x,y\in Q\in{\mathcal{D}}_{-a}$, then $x-y\in R.$ By Jensen’s inequality ($p\geq 1$), Hölder’s inequality ($r/p\geq 1$), and (20), we obtain
$\|\vec{f}-E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}$
$=\left(\sum_{Q\in{\mathcal{D}}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|\frac{1}{|Q_{-a,0}|}\int_{Q_{-a,0}}W^{1/p}(x)\left(\vec{f}(x)-\chi_{Q_{-a,0}}(x-x_{Q})\vec{f}(y-x_{Q})\right){\mathrm{d}}y\right|^{p}{\mathrm{d}}x\right)^{r/p}\right)^{1/r}$
$\leq\left(\sum_{Q\in{\mathcal{D}}}W(Q)^{r/t-r/p}\left(\frac{1}{|Q_{-a,0}|}\int_{Q_{-a,0}}\int_{Q}\left|W^{1/p}(x)\left(\vec{f}(x)-\chi_{Q_{-a,0}}(x-x_{Q})\vec{f}(y-x_{Q})\right)\right|^{p}{\mathrm{d}}x{\mathrm{d}}y\right)^{r/p}\right)^{1/r}$
$\leq\left(\sum_{Q\in{\mathcal{D}}}W(Q)^{r/t-r/p}\frac{1}{|Q_{-a,0}|}\int_{Q_{-a,0}}\left(\int_{Q}\left|W^{1/p}(x)\left(\vec{f}(x)-\chi_{Q_{-a,0}}(x-x_{Q})\vec{f}(y-x_{Q})\right)\right|^{p}{\mathrm{d}}x\right)^{r/p}{\mathrm{d}}y\right)^{1/r}$
$\leq\frac{1}{|Q_{-a,0}|^{1/r}}|Q_{-a,0}|^{1/r}\sup_{y\in R}\left(\sum_{Q\in{\mathcal{D}}}W(Q)^{r/t-r/p}\left(\int_{Q}\left|W^{1/p}(x)(\vec{f}(x)-\vec{f}(x-y))\right|^{p}{\mathrm{d}}x\right)^{r/p}\right)^{1/r}\leq\epsilon,$
modified when $r=\infty$. Hence we prove that (8) is stronger than (18).
@@
where $E_{d,a}$ is same as (19).
Proof.
Assume that $\mathcal{F}\subset M_{p}^{t,r}(W)$ satisfies (i)-(iii). Given $\epsilon>0$ small enough, to prove the total boundedness of $\mathcal{F}$, it suffices to find a finite $\epsilon$-net of $\mathcal{F}$. Denote by $R_{i}:=[-2^{i},2^{i})^{n}$ for $i\in\mathbb{Z}$. Then from condition (ii), there exists a positive integer $m$ large enough such that
$\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-\vec{f}\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}<\epsilon.$ (21)
@@
$\sup_{\vec{f}\in\mathcal{F}}\|\vec{f}-E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}<\epsilon.$ (22)
There exists a sequence $\{Q_{j}\}_{j=1}^{N}$ of disjoint cubes in ${\mathcal{D}}_{-a}$ such that $R_{m}=\bigcup_{i=1}^{N}Q_{j}$, where $N=2^{(m+1-a)n}$. For any $\vec{f}\in\mathcal{F}$ and $x\in{\mathbb{R}^{n}}$, let
$\Phi(\vec{f})(x):=\begin{cases}\vec{f}_{Q_{j}}:=\frac{1}{|Q_{j}|}\int_{Q_{j}}\vec{f}(y){\mathrm{d}}y,&x\in Q_{j},j=1,2,\ldots,N,\\ \vec{0},&\operatorname{otherwise}.\end{cases}$
@@
$\Phi(\vec{f})(x)=E_{d,a}\vec{f}(x).$
Hence via (21) and (22), we have
$\|\vec{f}-\Phi(\vec{f})\|_{M_{p}^{t,r}(W)}$ $\leq\|(\vec{f}-\Phi(\vec{f}))\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}+\|(\vec{f}-\Phi(\vec{f}))\chi_{R_{m}^{c}}\|_{M_{p}^{t,r}(W)}$
$=\|(\vec{f}-E_{d,a}\vec{f})\chi_{R_{m}}\|_{M_{p}^{t,r}(W)}+\|\vec{f}\chi_{R_{m}^{c}}\|_{M_{p}^{t,r}(W)}$
$\leq\|(\vec{f}-E_{d,a}\vec{f})\|_{M_{p}^{t,r}(W)}+\epsilon\leq 2\epsilon.$ (23)
From (23), it suffices to show that $\Phi(\mathcal{F})$ is totally bounded in $M_{p}^{t,r}(W)$. And we have proved that $\Phi(\mathcal{F})$ is totally bounded in $M_{p}^{t,r}(W)$ in Theorem 3.2. Thus the proof of Theorem 4.1 is complete. ∎
Lemma 4.2 (Lemma 4.5, [10]).
@@
Proof.
First, for any $\vec{f}=(f_{1},f_{2},\ldots,f_{d})^{T}\in M_{p}^{t,r}(W)$, it suffices to show that $f_{i}\in L^{1}_{\operatorname{loc}}({\mathbb{R}^{n}})$ for each $i=1,2,\ldots,d.$ Since $W\in\mathcal{A}_{p}$, then by (3)
@@
$\int_{Q}|\vec{f}(x)|{\mathrm{d}}x$ $=\int_{Q}|\vec{f}(x)|\|W^{-1}(x)\|^{-1/p}\|W^{-1}(x)\|^{1/p}{\mathrm{d}}x$
$\leq\left(\int_{Q}|\vec{f}(x)|^{p}\|W^{-1}(x)\|^{-1}{\mathrm{d}}x\right)^{1/p}\left(\int_{Q}\|W^{-1}(x)\|^{p^{\prime}/p}{\mathrm{d}}x\right)^{1/p^{\prime}}$
$\leq C_{Q,W,p}\left(\int_{Q}|W^{1/p}(x)\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{1/p}\lesssim\|\vec{f}\|_{M_{p}^{t,r}(W)}<\infty.$
@@
Remark 4.5.
In the sequel, let $p,p^{\prime},W,\tilde{d},\widetilde{W},\tilde{\tilde{d}}$ have the same meaning as in Lemma 2.9. Let
$\tilde{\beta}:=\begin{cases}n,&\operatorname{if}\;p=1,\\ n+\tilde{\tilde{d}}p/p^{\prime},&\operatorname{if}\;1<p<\infty.\end{cases}$ (24)
Then we claim that for $j+a\in\mathbb{N}_{0}$,
$W((j+a)_{\operatorname{pa}}Q)\lesssim\begin{cases}2^{(j+a)n}W(Q),&\operatorname{if}p=1,\\ 2^{(j+a)(n+\tilde{\tilde{d}}p/p^{\prime})}W(Q),&\operatorname{if}1<p<\infty.\end{cases}$
Indeed, when $p=1$, by Lemma 2.9, we have
$W((j+a)_{\operatorname{pa}}Q)\approx|(j+a)_{\operatorname{pa}}Q|\|A_{(j+a)_{\operatorname{pa}}Q}\|\lesssim 2^{(j+a)n}|Q|\|A_{Q}\|\approx 2^{(j+a)n}W(Q).$
When $p\in(1,\infty)$, by Lemma 2.9, we obtain
$W((j+a)_{\operatorname{pa}}Q)$ $\approx|(j+a)_{\operatorname{pa}}Q|\|A_{(j+a)_{\operatorname{pa}}Q}\|^{p}\lesssim 2^{(j+a)n}|Q|2^{(j+a)\tilde{\tilde{d}}p/p^{\prime}}\|A_{Q}\|^{p}\approx 2^{(j+a)(n+\tilde{\tilde{d}}p/p^{\prime})}W(Q).$
Thus we prove the claim. Now we show this estimate is better than $W((j+a)_{\operatorname{pa}}Q)\lesssim 2^{(j+a)\beta}W(Q)$. Indeed, when $p=1$, it is obvious since $\beta\geq n$. If $p>1$, from [3, Lemma 2.11 and Corollary 2.16], for $i\in\mathbb{N}_{0}$, we deduce that
$\|A_{i_{\operatorname{pa}}Q}A_{Q}^{-1}\|^{p}$ $\approx\frac{1}{|i_{\operatorname{pa}}Q|}\int_{i_{\operatorname{pa}}Q}\left\|W^{1/p}(x)A_{Q}^{-1}\right\|^{p}{\mathrm{d}}x$
$\approx\frac{1}{|i_{\operatorname{pa}}Q|}\int_{i_{\operatorname{pa}}Q}\left(\frac{1}{|Q|}\int_{Q}\left\|W^{1/p}(x)W^{-1/p}(y)\right\|^{p^{\prime}}{\mathrm{d}}y\right)^{p/p^{\prime}}{\mathrm{d}}x.$
From [12, Lemma 2.2], we have $\|A_{i_{\operatorname{pa}}Q}A_{Q}^{-1}\|^{p}\lesssim 2^{i(\beta-n)}$. Hence
$\frac{1}{|i_{\operatorname{pa}}Q|}\int_{i_{\operatorname{pa}}Q}\left(\frac{1}{|Q|}\int_{Q}\left\|W^{1/p}(x)W^{-1/p}(y)\right\|^{p^{\prime}}{\mathrm{d}}y\right)^{p/p^{\prime}}{\mathrm{d}}x\lesssim 2^{i(\beta-n)}.$
This, together with [3, Proposition 2.28(ii)], further implies that $W^{-1/(p-1)}\in\mathcal{A}_{p^{\prime}}$ has the $\mathcal{A}_{p^{\prime}}$-dimension $\tilde{\tilde{d}}:=(\beta-n)/(p-1)$ and hence
$n+\tilde{\tilde{d}}p/p^{\prime}=\beta.$
Hence we show this estimate is better than $W((j+a)_{\operatorname{pa}}Q)\lesssim 2^{(j+a)\beta}W(Q)$.
Theorem 4.6.
Let $W\in\mathcal{A}_{p}$ has the $\mathcal{A}_{p}$-dimension $\tilde{d}\in[0,n)$. Let $\tilde{\beta}$ be the same with (24).
(i) Let $1\leq p<t<r<\infty$ . If $nr/p+\tilde{d}r/p+n-\tilde{\beta}r(1/t-1/p)<0$, then for each $a\in\mathbb{Z}$, $E_{d,a}$ is a bounded operator on matrix weighted Bourgain-Morrey spaces $M_{p}^{t,r}(W)$.
(ii) Let $1\leq p\leq t<r=\infty$ . If $\tilde{d}/p-n/p-\tilde{\beta}(1/t-1/p)\leq 0$, then for each $a\in\mathbb{Z}$, $E_{d,a}$ is a bounded operator on matrix weighted Bourgain-Morrey spaces $M_{p}^{t,r}(W)$.
Proof.
@@
$+\left(\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}\left(\int_{Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}\right)^{1/r}=:S_{1}+S_{2}.$
@@
$S_{1}$ $\leq C_{n,d,p,W}\left(\sum_{j\leq-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}\left(\int_{Q}|W^{1/p}(x)\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}\right)^{1/r}\lesssim\|\vec{f}\|_{M_{p}^{t,r}(W)}.$
When j>aj>-a, we denote by $(j+a)_{\operatorname{pa}}Q\in{\mathcal{D}}_{-a}$ the $(j+a)$-th dyadic parent of $Q\in{\mathcal{D}}_{j}$. Let $\{A_{Q}\}_{Q\in{\mathcal{D}}}$ be the reducing operator of order $p$ for $W$. By Lemmas 2.9, we have
$\|A_{Q}A_{(j+a)_{\operatorname{pa}}Q}^{-1}\|\leq c2^{(j+a)\tilde{d}/p}.$ (25)
In $S_{2}$, using (5) and (25) , we have
@@
$=\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}\left(\frac{|Q|}{|Q|}\int_{Q}\left|W^{1/p}(x)\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}\vec{f}(y){\mathrm{d}}y\right|^{p}{\mathrm{d}}x\right)^{r/p}$
@@
$\lesssim\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}|Q|^{r/p}\left(2^{(j+a)\tilde{d}/p}\left|A_{(j+a)_{\operatorname{pa}}Q}\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}\vec{f}(y){\mathrm{d}}y\right|\right)^{r}$
$\approx\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}|Q|^{r/p}2^{(j+a)\tilde{d}r/p}\left(\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}\vec{f}(y){\mathrm{d}}y|^{p}{\mathrm{d}}x\right)^{r/p}$
$=\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}\left(\frac{1}{|(j+a)_{\operatorname{pa}}Q|}\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}$
$=\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}2^{-anr/p}\left(\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}.$
Remark that given $S\in{\mathcal{D}}_{-a}$, there are $2^{(j+a)n}$ cubes $R$ such that $(j+a)_{\operatorname{pa}}R=S$. From Remark 4.5 and 1/t1/p<01/t-1/p<0, we have
$W(Q)^{1/t-1/p}\lesssim\begin{cases}2^{-(j+a)n(1/t-1)}W((j+a)_{\operatorname{pa}}Q)^{1/t-1},&\operatorname{if}p=1,\\ 2^{-(j+a)(n+\tilde{\tilde{d}}p/p^{\prime})(1/t-1/p)}W((j+a)_{\operatorname{pa}}Q)^{1/t-1/p},&\operatorname{if}1<p<\infty.\end{cases}$
That is $W(Q)^{1/t-1/p}\lesssim 2^{-(j+a)\tilde{\beta}(1/t-1/p)}W((j+a)_{\operatorname{pa}}Q)^{1/t-1/p}$. Hence, by Lemma 4.4 and $nr/p+\tilde{d}r/p+n-\tilde{\beta}r(1/t-1/p)<0$, we have
$\sum_{j>-a}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}2^{-anr/p}\left(\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}$
$=\sum_{j>-a}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}2^{-anr/p}\sum_{Q\in\mathcal{D}_{j}}W(Q)^{r/t-r/p}\left(\int_{(j+a)_{\operatorname{pa}}Q}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}$
$\leq\sum_{j>-a}2^{-jnr/p}2^{(j+a)\tilde{d}r/p}2^{-anr/p}2^{(j+a)n}2^{-(j+a)\tilde{\beta}(r/t-r/p)}\sum_{S\in\mathcal{D}_{-a}}W(S)^{r/t-r/p}\left(\int_{S}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{r/p}$
$\lesssim\|\vec{f}\|_{M_{p}^{t,r}(W)}^{r}.$
@@
$W(Q_{j,k})^{1/t-1/p}\left(\int_{Q_{j,k}}|W^{1/p}(x)E_{d,a}\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{1/p}\lesssim W(Q_{j,k})^{1/t-1/p}\left(\int_{Q_{j,k}}|W^{1/p}(x)\vec{f}(x)|^{p}{\mathrm{d}}x\right)^{1/p}.$
If j>aj>-a, we denote by $(j+a)_{\operatorname{pa}}Q_{j,k}\in{\mathcal{D}}_{-a}$ the $(j+a)$-th dyadic parent of $Q_{j,k}\in{\mathcal{D}}_{j}$.